Problem 4.20 Constants | Periodic Table Part A f 3.585 g of ethanol C2 H5 OH(1) is burned completely in a bomb calorimeter at 298.15 K, the heat produced is 106.0 kJ Calculate A Howmbin for ethanol at 298.15 K. Calculate A Hcom Express your answer to four significant figures and include the appropriate units A Hcombuntion-282.4 kJ mol combustiorn Previous Answers Request Answer X Incorrect; Try Again; 4 attempts remaining
Answer #1

Molar mass of C2H5OH,

MM = 2*MM(C) + 6*MM(H) + 1*MM(O)

= 2*12.01 + 6*1.008 + 1*16.0

= 46.068 g/mol

mass(C2H5OH)= 3.585 g

use:

number of mol of C2H5OH,

n = mass of C2H5OH/molar mass of C2H5OH

=(3.585 g)/(46.07 g/mol)

= 7.782*10^-2 mol

Q = -106.0 KJ

delta H = Q / number of mol

= -106.0 KJ / ( 7.782*10^-2 mol)

= -1362 KJ/mol

Answer: -1362 KJ/mol